OpenStudy (anonymous):

Help please Part 1: Solve the inequality 5x - 2<= 2x + 10 Part 2: Use complete sentences to describe the graph of the solution from Part 1. Be sure to include the type of circle and which direction the shading will be.

OpenStudy (texaschic101):

5x - 2 <= 2x + 10 first you need to get the x's on one side and everything else on the other side subtract 2x from both sides because what you do one one side, you have to do on the other side. 5x - 2x - 2 <= 2x - 2x + 10 (combine like terms) 3x - 2 <= 10 Now you need to add 2 to both sides to get 3x by itself 3x - 2 + 2 <= 10 + 2 (combine like terms) 3x <= 12 Now to get x by itself, you have to divide both sides by 3 3x/3 <= 12/3 x <= 4 Do you have any questions ?

OpenStudy (anonymous):

No.Thank you very much.

OpenStudy (texaschic101):

your welcome :)